OpenStudy (anonymous):

WILL MEDAL 3 - 10 = ? @TheSmartOne

OpenStudy (anonymous):

7 right?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

it is not 7

OpenStudy (alexandervonhumboldt2):

3-10=-(10-3)=?

OpenStudy (anonymous):

Yes if it was 10 - 3... I'm asking 3 - 10

OpenStudy (alexandervonhumboldt2):

yes i know. i say 3-10=-(10-3)=?

TheSmartOne (thesmartone):

Hint: \(\sf a - b = -b + a = \boxed{-1( b - a)}\)

TheSmartOne (thesmartone):

that hint is useful in this question because the second number is larger than the first :)

OpenStudy (anonymous):

I don't understand @TheSmartOne

OpenStudy (dbzfan836):

@Smart_Owl_Student what happens you lose all positive?

OpenStudy (anonymous):

wow this crazy for me

OpenStudy (anonymous):

Ok guys thx! (-7) is the answer!

OpenStudy (dbzfan836):

np

OpenStudy (anonymous):

And the medal goes to....

OpenStudy (anonymous):

(Drum Roll)

OpenStudy (anonymous):

 Number lines always help. If you move left 10 from 3, where do you arrive? -7.

OpenStudy (anonymous):

@TheSmartOne ! Thx every1

TheSmartOne (thesmartone):

No problem :)